# Privacy-Preserving Federated Learning against Malicious Clients Based on Verifiable Functional Encryption

### v3

@@ Appendix D Proof of Theorem 5
Completeness. Perfect completeness follows from the fact that for all valid witnesses, it holds that $y=\frac{\langle x_{i},x_{0}\rangle}{\langle x_{i},x_{i}\rangle}$ and
@@
$g^{\langle L,x_{0}\rangle}(u^{m})^{\tau^{\prime}_{\alpha}}$ $=g^{\langle x_{i}+k\alpha,x_{0}\rangle}u^{\tau_{0}y_{i}+\tau^{\prime}_{1}\alpha}$
@@
$w^{L}(u^{m})^{\omega}$ $=w^{x_{i}+k\alpha}u^{s_{i}+\alpha\rho}$
$=w^{x_{i}}u^{s_{i}}(w^{k}u^{\rho})^{\alpha}=V_{i}(K)^{\alpha}$
@@
$L_{1}=x_{i}+k\alpha_{1}$ $\text{ and }L_{2}=x_{i}+k\alpha_{2}$
@@
$x_{i}$ $=L_{1}-k\alpha_{1}=L_{1}-\alpha_{1}\frac{L_{1}-L_{2}}{\alpha_{1}-\alpha_{2}}$
@@
to extract $x_{i},s_{i},\tau_{0},\tau_{1},\tau_{2},\tau^{\prime}_{1}$, and then further computation of
$t_{0}=\langle x_{i},x_{i}\rangle,\ t_{1}=2\langle k,x_{i}\rangle,\ t_{2}=\langle k,k\rangle,\ t^{\prime}_{1}=\langle k,x_{0}\rangle$
@@
Zero knowledge. To demonstrate the protocol achieves perfect honest-verifier zero-knowledge, we construct a simulator that, given a statement $\left(V_{i}\in\mathbb{G}_{1},y_{i}\in\mathbb{Z}_{p},\mathrm{com}_{i}\in\mathbb{G}^{2}_{1}\right)$, produces a simulated proof whose distribution is indistinguishable from that of a real proof generated through interaction between an honest prover and an honest verifier. All elements chosen by the simulator are either independently and uniformly distributed or fully determined by the verification equations. The simulator randomly chooses $T_{0},T_{2},\tau_{\alpha},\tau^{\prime}_{\alpha},\omega$ and $L$. Then, for the challenge $\alpha$ and the public value $x_{0}$, the values $t,T_{1},T^{\prime}_{1},K$, and $\mathrm{com}^{*}$ are computed according to the verification equations.
@@
$T^{\prime}_{1}$ $=\left(g^{\langle L,x_{0}\rangle}(u^{m})^{\tau^{\prime}_{\alpha}}T_{0}^{-y_{i}}\right)^{-\alpha}$
$K$ $=\left(w^{L}(u^{m})^{\omega}(V_{i})^{-1}\right)^{-\alpha}$
@@ Appendix E Lemma 5 and Proof
Let $lr$ denote the number of local training rounds. We define $w^{0}=W_{0}$, and after $lr$ rounds of local training on dataset $D_{0}$, the final model is denoted by $w^{lr}=W_{0}^{t}$. During the $i$-th round of training, the gradient descent update rule gives:
$w^{i+1}=w^{i}-\eta\nabla F(w^{i};D_{0}).$
@@
$\|w^{i+1}-W^{opt}\|^{2}$
$=$ $\|w^{i}-W^{opt}\|^{2}-2\eta\langle\nabla F(w^{i};D_{0}),w^{i}-W^{opt}\rangle$
$+\eta^{2}\|\nabla F(w^{i};D_{0})\|^{2}.$
@@
$F(w^{i})+\langle\nabla F(w^{i};D_{0}),(W^{opt}-w^{i})\rangle$
$+\frac{\mu}{2}\|W^{opt}-w^{i}\|^{2}$ $\leq F(W^{opt})$
$F(W^{opt})+\frac{\mu}{2}\|w^{i}-W^{opt}\|^{2}$ $\leq F(w^{i})$
$\|\nabla F(w^{i};D_{0})-\nabla F(W^{opt})\|^{2}$ $\leq L^{2}\|w^{i}-W^{opt}\|^{2}$
@@
$-2\eta\langle\nabla F(w^{i};D_{0}),w^{i}-W^{opt}\rangle$ $\leq-2\eta\mu\|w^{i}-W^{opt}\|^{2}$
$\eta^{2}\|\nabla F(w^{i};D_{0})\|^{2}$ $\leq\eta^{2}L\|w^{i}-W^{opt}\|^{2}$
@@
$\|w^{lr}-W^{opt}\|\leq\left(\sqrt{(1-2\eta\mu+\eta^{2}L^{2})}\right)^{lr}\|w^{0}-W^{opt}\|.$
Since $w^{0}=W_{0},w^{lr}=W_{0}^{t}$, the global model after $lr$ local training rounds satisfies:

### v4

@@ Appendix D Proof of Theorem 5
Completeness. Perfect completeness follows from the fact that for all valid witnesses, it holds that $y=\frac{\langle\vec{x}_{i},\vec{x}_{0}\rangle}{\langle\vec{x}_{i},\vec{x}_{i}\rangle}$ and
@@
$g^{\langle L,\vec{x}_{0}\rangle}(u^{m})^{\tau^{\prime}_{\alpha}}$ $=g^{\langle\vec{x}_{i}+k\alpha,\vec{x}_{0}\rangle}u^{\tau_{0}y_{i}+\tau^{\prime}_{1}\alpha}$
@@
$\vec{w}^{L}(u^{m})^{\omega}$ $=\vec{w}^{\vec{x}_{i}+k\alpha}u^{s_{i}+\alpha\rho}$
$=\vec{w}^{\vec{x}_{i}}u^{s_{i}}(\vec{w}^{k}u^{\rho})^{\alpha}=V_{i}(K)^{\alpha}$
@@
$L_{1}=\vec{x}_{i}+k\alpha_{1}$ $\text{ and }L_{2}=\vec{x}_{i}+k\alpha_{2}$
@@
$\vec{x}_{i}$ $=L_{1}-k\alpha_{1}=L_{1}-\alpha_{1}\frac{L_{1}-L_{2}}{\alpha_{1}-\alpha_{2}}$
@@
to extract $\vec{x}_{i},s_{i},\tau_{0},\tau_{1},\tau_{2},\tau^{\prime}_{1}$, and then further computation of
$t_{0}=\langle\vec{x}_{i},\vec{x}_{i}\rangle,\ t_{1}=2\langle k,\vec{x}_{i}\rangle,\ t_{2}=\langle k,k\rangle,\ t^{\prime}_{1}=\langle k,\vec{x}_{0}\rangle$
@@
Zero knowledge. To demonstrate the protocol achieves perfect honest-verifier zero-knowledge, we construct a simulator that, given a statement $\left(V_{i}\in\mathbb{G}_{1},y_{i}\in\mathbb{Z}_{p},\mathrm{com}_{i}\in\mathbb{G}^{2}_{1}\right)$, produces a simulated proof whose distribution is indistinguishable from that of a real proof generated through interaction between an honest prover and an honest verifier. All elements chosen by the simulator are either independently and uniformly distributed or fully determined by the verification equations. The simulator randomly chooses $T_{0},T_{2},\tau_{\alpha},\tau^{\prime}_{\alpha},\omega$ and $L$. Then, for the challenge $\alpha$ and the public value $\vec{x}_{0}$, the values $t,T_{1},T^{\prime}_{1},K$, and $\mathrm{com}^{*}$ are computed according to the verification equations.
@@
$T^{\prime}_{1}$ $=\left(g^{\langle L,\vec{x}_{0}\rangle}(u^{m})^{\tau^{\prime}_{\alpha}}T_{0}^{-y_{i}}\right)^{-\alpha}$
$K$ $=\left(\vec{w}^{L}(u^{m})^{\omega}(V_{i})^{-1}\right)^{-\alpha}$
@@ Appendix E Lemma 5 and Proof
Let $lr$ denote the number of local training rounds. We define $\vec{w}^{0}=W_{0}$, and after $lr$ rounds of local training on dataset $D_{0}$, the final model is denoted by $\vec{w}^{lr}=W_{0}^{t}$. During the $i$-th round of training, the gradient descent update rule gives:
$\vec{w}^{i+1}=\vec{w}^{i}-\eta\nabla F(\vec{w}^{i};D_{0}).$
@@
$\|\vec{w}^{i+1}-W^{opt}\|^{2}$
$=$ $\|\vec{w}^{i}-W^{opt}\|^{2}-2\eta\langle\nabla F(\vec{w}^{i};D_{0}),\vec{w}^{i}-W^{opt}\rangle$
$+\eta^{2}\|\nabla F(\vec{w}^{i};D_{0})\|^{2}.$
@@
$F(\vec{w}^{i})+\langle\nabla F(\vec{w}^{i};D_{0}),(W^{opt}-\vec{w}^{i})\rangle$
$+\frac{\mu}{2}\|W^{opt}-\vec{w}^{i}\|^{2}$ $\leq F(W^{opt})$
$F(W^{opt})+\frac{\mu}{2}\|\vec{w}^{i}-W^{opt}\|^{2}$ $\leq F(\vec{w}^{i})$
$\|\nabla F(\vec{w}^{i};D_{0})-\nabla F(W^{opt})\|^{2}$ $\leq L^{2}\|\vec{w}^{i}-W^{opt}\|^{2}$
@@
$-2\eta\langle\nabla F(\vec{w}^{i};D_{0}),\vec{w}^{i}-W^{opt}\rangle$ $\leq-2\eta\mu\|\vec{w}^{i}-W^{opt}\|^{2}$
$\eta^{2}\|\nabla F(\vec{w}^{i};D_{0})\|^{2}$ $\leq\eta^{2}L\|\vec{w}^{i}-W^{opt}\|^{2}$
@@
$\|\vec{w}^{lr}-W^{opt}\|\leq\left(\sqrt{(1-2\eta\mu+\eta^{2}L^{2})}\right)^{lr}\|\vec{w}^{0}-W^{opt}\|.$
Since $\vec{w}^{0}=W_{0},\vec{w}^{lr}=W_{0}^{t}$, the global model after $lr$ local training rounds satisfies: